OpenStudy (anonymous):

Evaluate. (16 + 32 ÷ 4)(21 – 6 • 3)

OpenStudy (xapproachesinfinity):

when you this sort of operation: you give priority the multiplication first then come division and later addiction and subtraction.

OpenStudy (anonymous):

@ManImStumped

OpenStudy (itsbribro):

use PEMDAS

OpenStudy (anonymous):

?

OpenStudy (itsbribro):

Perenthases Exponent Multiplication Division Addition Subtraction

OpenStudy (xapproachesinfinity):

is just order of operation what comes first as i explained it above but in you case you have exponent and you don't need to care about parenthesis

OpenStudy (xapproachesinfinity):

read my reply and try to do what i said okay

OpenStudy (itsbribro):

do you want me to work it out with you?

OpenStudy (itsbribro):

@m&m&m

OpenStudy (anonymous):

ya I don't get it

OpenStudy (itsbribro):

okay so your question is \[(16+32\div4)(21-6\times3)\]

OpenStudy (anonymous):

so 6*3 first right?

OpenStudy (itsbribro):

yes

OpenStudy (anonymous):

18

OpenStudy (itsbribro):

\[(16+32\div4)(21-18)\]

OpenStudy (itsbribro):

right

OpenStudy (itsbribro):

what next?

OpenStudy (anonymous):

21-8

OpenStudy (itsbribro):

nooo we would to multiplication and division first

OpenStudy (anonymous):

32/4?

OpenStudy (itsbribro):

yes

OpenStudy (itsbribro):

\[(16+8)(21-18)\]

OpenStudy (anonymous):

(24) (13)

OpenStudy (itsbribro):

21-18=

OpenStudy (itsbribro):

you have 21 candy bars and you eat 18 how many do you have left?

OpenStudy (anonymous):

16+8=24 21-18=13

OpenStudy (itsbribro):

16+8=24 yesss

OpenStudy (itsbribro):

21-18 does not =13

OpenStudy (anonymous):

oops 3

OpenStudy (itsbribro):

yeah u got it now you have (24)(3) you can drop the parenthases and add a multiplication sign 24 x 3

OpenStudy (itsbribro):

whats 24 x 3?

OpenStudy (anonymous):

72 thank you

OpenStudy (itsbribro):

your welcome :)

OpenStudy (itsbribro):

do you need anymore help?